Archive for October, 2009

Outage: Framestream Circuits

Posted: Thursday, October 29th, 2009 at 17:31 by Adam Heath

We are aware of an outage currently affecting a number of Framestream circuits.  It would appear the outage is localised (West Midlands) and our supplier has been informed.

Please accept our apologies for any loss to your service, more information will be provided as and when we receive it.

  1. Martyn Haynes says:

    Service has been restored. More to follow.

  2. Martyn Haynes says:

    Our supplier reports there is fault with one of the ATM links and have re-routed traffic via existing paths.

Outage: Stoke to Manchester

Posted: Thursday, October 29th, 2009 at 12:31 by David Derrick

Our link between Manchester and Stoke-on-Trent is currently down. Our supplier is aware of a problem and is investigating. All traffic is routing via alternate paths.

  1. David Derrick says:

    Our supplier has located the fibre break in the Kidsgrove area and engineers are heading to the site. No ETR at present.

  2. David Derrick says:

    The cable break is repaired and all engineering work is closed off. Traffic is back on its normal paths.

Incoming VoIP Calls To Ported In Numbers

Posted: Wednesday, October 28th, 2009 at 08:55 by Adam Heath

We are currently experiencing problems with incoming VoIP calls where the number has been ported from a different supplier to us. The problem is with one of our suppliers and engineers are investigating. Callers are hearing a recorded announcement.

  1. Jon Farmer says:

    We have been advised this fault is now fixed.

Unexplained session drops

Posted: Wednesday, October 21st, 2009 at 22:47 by Steve Lalonde

We are investigating a problem with xDSL sessions dropping from multiple LNS in the Telehouse cluster.

  1. Steve Lalonde says:

    The problem seems to have disappeared and the LNS cluster in Telehouse is stable again.

telehouse-east2.core

Posted: Tuesday, October 20th, 2009 at 17:26 by Steve Lalonde

We are investigating an issue with 1 of the blades in the above router.

This only affects customers connected to the affected blade.

  1. Steve Lalonde says:

    The blade has failed, We have dispatched an engineer to site with a replacement blade.

  2. David Derrick says:

    The blade has been replaced and service is now restored. Customers who are still seeing a problem with their service in Telehouse East are asked to check their side of the connection and contact support if the service is still down.

Emergency Maintenance: VoIP Platform

Posted: Friday, October 16th, 2009 at 12:58 by Adam Heath

Tuesday 20th October 2009, 04:00hrs to 05:00hrs.

It is necessary to perform emergency work on the database servers controlling our VoIP platform. It is expected that this work will take between 30 and 45 minutes to complete. It is anticipated that during this window customers may experience loss of inbound and outbound calls, this will include calls internal to the Entanet VoIP platform.

Additionally the Entanet Support Helpdesk will be unavailable during this period.

We apologise for the inconvenience that this will cause and wish to reassure you that any downtime will be kept to a minimum.

  1. Jon Farmer says:

    This work is now completed and full VoIP service has been resumed.

Outage: Bristol 21CN Issues

Posted: Friday, October 16th, 2009 at 10:01 by Adam Heath

We are currently experiencing an authentication issue with connections into our 21CN Bristol node on the DSL platform. As a consequence users on this node may not be able to connect at present. This has been reported to BT who have advised they are looking into the issue as multiple service providers have reported similar issues in this area.

  1. Marc Woolfenden says:

    We have been advised by BT that the current situation is believed to be linked to a BRAS issue within their network earlier this morning and are continuing to investigate.

    We will advise of further updates as they become available.

  2. Marc Woolfenden says:

    We have been advised by BT that following a restart of some BRAS services at 13:00 connectivity should now be restored to all users.

    If any users continue to experience problems we would recommend reporting them via the normal Support channels.

DSL Authentication Issues

Posted: Thursday, October 15th, 2009 at 17:29 by Neil Watson

We are currently aware of an issue that is affecting authentication for DSL users. As a consequence users may not be able to connect at present. Our engineers are working to correct the issue as a matter of urgency and we will update with more info as soon as possible.

  1. Neil Watson says:

    Services have now returned to normal, we are currently checking for the cause as nothing on our network indicates where the problem originated. We are working with our suppliers to see where the problem may be and will continue to update.

    If any users are still unable to authenticate, please power down the equipment for 5 – 10 minutes to allow the previous cached session to expire.

Outage: Manchester to London 10 Gigabit

Posted: Thursday, October 15th, 2009 at 08:04 by Steve Lalonde

Our main Manchester to London circuit is down, a ticket has been raised with the carrier.

Traffic usually on this route is using an alternative path, there may be a small increase in latency due to the change in routing.

  1. Steve Lalonde says:

    Initial investigations point to a fibre problem in our carriers ODF in Telehouse North. The carrier is working with Telehouse remote hands to rectify the fault.

  2. Steve Lalonde says:

    Service has been restored and traffic is now running across the preferred paths.

Increased Latency to Peterborough

Posted: Wednesday, October 14th, 2009 at 17:23 by Adam Heath

Currently the circuit between London and Peterborough is experiencing some packet loss. Traffic has been re-routed to alternative paths to avoid this link which means increased latency to Peterborough may be present. A fault has been logged with our Carrier and further updates will follow as soon as we have more information.

  1. Steve Lalonde says:

    The packet loss problem has now been resolved and traffic is now back on the preferred path.